OpenStudy (twizttiez):

How many solutions does this system of equations have? A. none B. exactly one C. exactly two D. infinitely many

OpenStudy (twizttiez):

@Write0The0Right0For0Wrong0Doing

OpenStudy (anonymous):

i say b.. but u can question that.. bc the answers r y=x/2+3 and y=x/2+0

OpenStudy (twizttiez):

Well I'm going to trust you :D

OpenStudy (anonymous):

lol.. what grade math is this?

OpenStudy (twizttiez):

Just a moment only a couple more

OpenStudy (anonymous):

ok

OpenStudy (twizttiez):

Not very high i suck at math English is my strong point

OpenStudy (anonymous):

lol.. what school u in?

OpenStudy (twizttiez):

English and all other subjects I'm in homeschooling at the moment but i would be going to Trinity christian academy or some other private school my parents just didn't sign me up quick enough

OpenStudy (anonymous):

lol..i am in flvs:)

OpenStudy (twizttiez):

K12 is what I'm in and i hate it all the kids in these stupid asf class connect Christ they're annoying xD

OpenStudy (anonymous):

lol...yea i bet well i am in 10grade